( a ) At t : 0, deposit of 5,000 is made in a saving account that pays 6% annually. Each year the deposit is increased by 200. How much money is available to the depositor at the end of 60 years? (b). An X-Ray machine at a dental office is classified as a MACRS 7-year property. It costs 12,000 and has an expected life of 12 years. The salvage value at the end of 12 years is expected to be 4,000. Using MACRS depreciation what is the book value at the end of 3 years
Answer #1

1.PV=$5000

Rate=6%

Increase = $200 each year

N=60

Amount = 5000*1.06^60 + 200*(P/G,6%,60)*1.06^60

=164938.5+ 200*239.043*1.06^60

=1,742,034

2.Cost = 12000

Less Depreciation

Year 1= 14.29%

Year 2= 24.49%

Year 3= 17.49%

Ending book value =12000-56.27%*12000 =

5247.6